OpenStudy (spid):

What is 6x-4b-73d+62j

OpenStudy (spid):

So do I minus 6-4

OpenStudy (spid):

Or do I add

OpenStudy (anonymous):

\[x=\frac{ 1 }{ 6 }(4b+73d-62j)\]

OpenStudy (spid):

I don't understand this problem

OpenStudy (anonymous):

are you solving for x?

OpenStudy (anonymous):

6x-4b-73d+62j cannot be simplified without additional information.

OpenStudy (spid):

So do I minus 6-4 and I am solving x

OpenStudy (anonymous):

your answer is above if you are solving for x

OpenStudy (spid):

Ok so 6-4 is 2 so I minus 73 to 62 right

OpenStudy (anonymous):

6x-4b-73d+62j is not an equation. Solving has no meaning.

OpenStudy (anonymous):

6x-4b-73d+62j=0 6x = 4b+73d-62j x=1/6(4b+73d-62j)

OpenStudy (spid):

Oh I see now

OpenStudy (spid):

Ok I got another question

OpenStudy (anonymous):

@robtobey he is solving for x

OpenStudy (anonymous):

Where is the equal sign in the original expression?

OpenStudy (anonymous):

i am assuming their is one which he didn't show

OpenStudy (spid):

@krypton

OpenStudy (spid):

@robtobey

OpenStudy (anonymous):

what?

OpenStudy (spid):

I'm sorry the equal sign is 9 I forgot to mention it